I am 61 and my niece is 6. Given this information, how do I write the following alegrabic equation? How old in years will each of us be when I am 3 times as old as my niece?
Answer by unlockmath(1688)

Hello,
Here we go. First set up x as the number of years. So the (61+x) will be you and your niece will be (6+x). Since you'll be 3 times older we can set her age as
3(6+x). Both of these will then equal each other so the equation is:
61+x=3(6+x) Rewritten it will be:
61+x=18+3x Subtract x from both sides gives us:
61=18+2x Subtract 18 from both sides to be:
43=2x Now divide both sides by 2
x=21.5 So now we add this to 6 which will be the niece's age:
27.5 years old and you are
82.5 years old
Make sense?
RJ
Check out a new math book at:
www.math-unlock.com